# Deformations of pseudocharacters and Mazur’s finiteness condition

### v1

@@ 1. Introduction
If $\Omega$ is any algebraically closed field of characteristic zero then a semisimple representation $\rho:\Gamma\rightarrow\operatorname{GL}_{n}(\Omega)$ of a group $\Gamma$ is determined by its trace. Anyone who worked out a character table of a finite group can appreciate that this is a useful result. Conversely, one might wonder when a conjugation invariant function from $\Gamma$ to $\Omega$ is a trace of a representation of $\Gamma$. Lafforgue’s pseudocharacters, whose definition we recall in Section 2, provide an answer to these questions for any (generalised) reductive group scheme $G$ instead of $\operatorname{GL}_{n}$ and without restriction on the characteristic of $\Omega$. They were introduced by V. Lafforgue in his work [11] on the Langlands correspondence in the function field case in order to associate Galois representations to automorphic forms. Since then the theory has been developed further by Böckle–Harris–Khare–Thorne [2], Dat–Helm–Kurinczuk–Moss [7], Emerson–Morel [9], Weidner [18], Zhu [19], and J.Q. [14].
@@ 1.1. Results
where $K(p)$ is the maximal pro-$p$ quotient of $K$. We may view $\overline{\Theta}$ as a $G$-pseudocharacter of $Q$. Moreover, any continuous pseudocharacter of $Q$ becomes a continuous pseudocharacter of $\Gamma$ after composing it with the quotient map. Hence, $\operatorname{Def}^{Q}_{\overline{\Theta}}$ is a closed subfunctor of $\operatorname{Def}^{\Gamma}_{\overline{\Theta}}$, which induces a surjection $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}$.
@@
The map $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}$ is an isomorphism.
A profinite group $\Gamma$ satisfies Mazur’s $p$-finiteness condition $\Phi_{p}$ if for every open subgroup $\Gamma^{\prime}$ of $\Gamma$ the space of continuous group homomorphisms $\operatorname{Hom}^{\mathrm{cont}}(\Gamma^{\prime},\mathbb{F}_{p})$ is finite. The Hermite–Minkowski theorem implies that the Galois group $\operatorname{Gal}(F_{S}/F)$, where $F$ is a number field and $F_{S}$ is a maximal extension of $F$ unramified outside a finite set of primes $S$ of $F$, satisfies $\Phi_{p}$. It is not known if these groups are topologically finitely generated. The deformation rings of representations of $\operatorname{Gal}(F_{S}/F)$ are frequently encountered in arithmetic applications and $\Gamma=\operatorname{Gal}(F_{S}/F)$ is the main example that we have in mind in this paper.
If $\Gamma$ satisfies $\Phi_{p}$ then it is easy to see that $K(p)$ is a finitely generated pro-$p$ subgroup and hence $Q$ is a finitely generated profinite group. It was shown in [14] that the ring $R^{\mathrm{ps},Q}_{\overline{\Theta}}$ is noetherian and hence we obtain:
@@
If $\Gamma$ satisfies $\Phi_{p}$ then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}=X^{\mathrm{gen},Q}_{\overline{\Theta}}$.
@@
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$ is represented by a finite type $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$-algebra.
@@
If $\Gamma$ satisfies $\Phi_{p}$ then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}\sslash G^{0}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ is a finite adequate homeomorphism in the sense of [1, Definition 3.3.1].
@@
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then the functor $\tilde{X}_{G}$ is representable by a quasi-Stein rigid analytic space $X_{G}$. The $\overline{L}$-points of $X_{G}$ are in canonical bijection with $G^{0}(\overline{L})$-conjugacy classes of continuous $G$-semisimple representations $\rho:\Gamma\rightarrow G(\overline{L})$.
@@ 1.2. What was known and tried before
Let us recall the previous related results. If $G=\operatorname{GL}_{n}$ then using the work of Emerson–Morel [9] one can work with Chenevier’s determinant laws instead of $\operatorname{GL}_{n}$-pseudocharacters. In this context 1.1 is proven by Chenevier in [5]: as part of the theory one gets a Cayley–Hamilton algebra $E$ together with a continuous representation $\rho^{\mathrm{univ}}:\Gamma\rightarrow E^{\times}$. One may then show that $\rho^{\mathrm{univ}}$ factors through $Q$ and deduce 1.1 from this, see [5, Lemma 3.8] and [12, Lemma 5.5] for details.
If $G$ is any generalised reductive group then the Cayley–Hamilton algebra is not available anymore. However, one could still try to deduce 1.2 from the $\operatorname{GL}_{n}$-case by choosing a closed immersion $\tau:G\hookrightarrow\operatorname{GL}_{n}$ and study the map $R^{\mathrm{ps},\Gamma}_{\tau\circ\overline{\Theta}}\rightarrow R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ induced by composition with $\tau$. If the map is finite then one may deduce from the $\operatorname{GL}_{n}$-case that $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ is noetherian. This approach was carried out by J.Q. in [14, Section 5.3] for classical groups under the assumption $p>2$ in the orthogonal cases. However, this strategy runs into difficult invariant theory questions in general.
If $\Gamma$ is a finitely generated group then J.Q. proves 1.2 by showing that the tangent space is finite dimensional, which is equivalent to showing that $\operatorname{Def}_{\overline{\Theta}}^{\Gamma}(k[\varepsilon])$ is finite. Let us informally explain the underlying difficulty if we drop the assumption that $\Gamma$ is finitely generated. The $G$-pseudocharacter $\Theta$ valued in the ring of dual numbers $k[\varepsilon]$ over $k$ consists of $\mathcal{O}$-algebra homomorphisms $\Theta_{n}:\mathcal{O}[G^{n}]^{G^{0}}\rightarrow\mathcal{C}(\Gamma^{n},k[\varepsilon])$ for $n\geq 1$ satisfying certain compatibility relations, where $\mathcal{C}(\Gamma^{n},k[\varepsilon])$ is the set of continuous maps. Proving finiteness assertions such as finite dimensionality of the tangent space amounts to showing that finitely many $\Theta_{n}$ determine $\Theta$ uniquely. This is feasible when the group is finitely generated, but the argument runs into trouble when this assumption is dropped. A similar problem arises in 1.6. One has to show that for an affinoid algebra $A$ a continuous $A$-valued $G$-pseudocharacter $\Theta$ takes values in some formal model of $A$. It is easy to show that each $\Theta_{n}$ takes values in $\mathcal{C}(\Gamma^{n},\mathcal{A}_{n})$ for some formal model $\mathcal{A}_{n}$ of $A$, but it is not clear that there is a formal model that works for all $\Theta_{n}$.
@@ 1.3. What we actually do
Let us explain the proof of 1.1. We first prove 1.1 for finitely generated free profinite groups $\Gamma$. We show that the map between deformation rings $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$, induced by mapping a deformation of $\overline{\rho}$ to its $G$-pseudocharacter, is injective. This key step is carried out in Section 4. Its proof uses the fact that one may describe the moduli space of $G$-pseudocharacters of a discrete finitely generated free group $F$ as a GIT-quotient of a moduli space of representations of $F$ into $G$. We show that if $\Gamma$ is the profinite completion of $F$ then the rings $R^{\mathrm{ps}}_{\overline{\Theta}}$ and $R^{\square}_{\overline{\rho}}$ are naturally isomorphic to completions of the local rings at points $\Theta|_{F}$ and $\overline{\rho}|_{F}$ in the respective moduli spaces. The assertion about injectivity boils down to commutative algebra, which is carried out in Section 3. We note that if we drop the assumption that $\Gamma$ is a free profinite group then it is not known whether $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ is injective, even when $G=\operatorname{GL}_{n}$, see [5, Remark 1.25].
@@
1.4. Applications
@@ 1.5. Acknowledgements
V.P. thanks Ulrich Görtz and Marc Levine for a stimulating discussion regarding Section 3. The authors thank Toby Gee for his comments.
@@ 2. Lafforgue’s pseudocharacters
The functor $A\mapsto\mathrm{PC}^{\Gamma}_{G}(A)$ is representable by an $\mathcal{O}$-algebra [14, Theorem 3.19]. Let us write $\operatorname{Rep}^{\Gamma}_{G}(A):=\operatorname{Hom}(\Gamma,G(A))$. The functor $A\mapsto\operatorname{Rep}^{\Gamma}_{G}(A)$ is also representable by an $\mathcal{O}$-algebra. To a representation $\rho:\Gamma\to G(A)$, we associate a $G$-pseudocharacter $\Theta_{\rho}\in\mathrm{PC}^{\Gamma}_{G}(A)$ by the formula
@@ 2.1. Continuous pseudocharacters
If $\Gamma$ is a topological group, $A$ is a topological ring and for all $n\geq 1$, the map $\Theta_{n}$ has image in the set $\mathcal{C}(\Gamma^{n},A)$ of continuous maps $\Gamma^{n}\to A$, then we say that $\Theta$ is continuous. We denote the set of continuous $G$-pseudocharacters of $\Gamma$ with values in $A$ by $\mathrm{cPC}_{G}^{\Gamma}(A)$.
@@ 2.2. Kernels
It is shown in [14, Lemma 3.11] that $\operatorname{Ker}(\Theta)$ is a normal subgroup of $\Gamma$.
@@ 2.3. Deformations of $G$-pseudocharacters
Let $\mathcal{O}$ be the ring of integers in a finite extension $L$ of $\mathbb{Q}_{p}$ with residue field $k$. Let $\Gamma$ be a profinite group and let $\overline{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(k)$. We recall a deformation problem of $\overline{\Theta}$ introduced in [14, Section 5]. Let $\mathfrak{A}_{\mathcal{O}}$ be the category of artinian local $\mathcal{O}$-algebras with residue field $k$. We define the deformation functor of $\overline{\Theta}$
@@
that sends an object $A\in\mathfrak{A}_{\mathcal{O}}$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\Gamma$ over $A$ with $\Theta\otimes_{A}k=\overline{\Theta}$. It is pro-representable by a local profinite $\mathcal{O}$-algebra $R^{\mathrm{ps}}_{\overline{\Theta}}$ with residue field $k$ by [14, Theorem 5.4].
@@
Denote by $\Theta^{u}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\overline{\Theta}})$ the universal $G$-pseudocharacter. If $A$ is an $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra we will write $\Theta^{u}_{|A}$ for the specialisation of $\Theta^{u}$ along $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A$. Since $R^{\mathrm{ps}}_{\overline{\Theta}}$ is profinite and hence Hausdorff, $\operatorname{Ker}(\Theta^{u})$ is a closed normal subgroup of $\Gamma$.

### v2

@@ 1. Introduction
If $\Omega$ is any algebraically closed field of characteristic zero then a semisimple representation $\rho:\Gamma\rightarrow\operatorname{GL}_{n}(\Omega)$ of a group $\Gamma$ is determined by its trace. Anyone who worked out a character table of a finite group can appreciate that this is a useful result. Conversely, one might wonder when a conjugation invariant function from $\Gamma$ to $\Omega$ is a trace of a representation of $\Gamma$. Lafforgue’s $G$-pseudocharacters, whose definition we recall in Section 2, provide an answer to these questions for any (generalised) reductive group scheme $G$ instead of $\operatorname{GL}_{n}$ and without restriction on the characteristic of $\Omega$. They were introduced by V. Lafforgue in his work [12] on the Langlands correspondence in the function field case in order to associate Galois representations to automorphic forms. Since then the theory has been developed further by Böckle–Harris–Khare–Thorne [3], Dat–Helm–Kurinczuk–Moss [8], Emerson–Morel [10], Weidner [21], Zhu [22], and J.Q. [16].
@@ 1.1. Results
where $K(p)$ is the maximal pro-$p$ quotient of $K$. We show in 5.3 that $Q$ does not depend on the choice of $\overline{\rho}$. We may view $\overline{\Theta}$ as a $G$-pseudocharacter of $Q$. Moreover, any continuous pseudocharacter of $Q$ becomes a continuous pseudocharacter of $\Gamma$ after composing it with the quotient map. Hence, $\operatorname{Def}^{Q}_{\overline{\Theta}}$ is a closed subfunctor of $\operatorname{Def}^{\Gamma}_{\overline{\Theta}}$, which induces a surjection $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}$. The following theorem is the technical heart of the paper.
@@
The map $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}\twoheadrightarrow R^{\mathrm{ps},Q}_{\overline{\Theta}}$ is an isomorphism. In particular, every deformation of $\overline{\Theta}$ factors through $Q$.
A profinite group $\Gamma$ satisfies Mazur’s $p$-finiteness condition $\Phi_{p}$ if for every open subgroup $\Gamma^{\prime}$ of $\Gamma$ the space of continuous group homomorphisms $\operatorname{Hom}^{\mathrm{cont}}(\Gamma^{\prime},\mathbb{F}_{p})$ is finite. If $\Gamma$ satisfies $\Phi_{p}$ then it is easy to see that $K(p)$ is a finitely generated pro-$p$ subgroup and hence $Q$ is a finitely generated [^1] Throughout the paper this means that the group in question is finitely generated as a topological group. profinite group. It was shown in [16] that the ring $R^{\mathrm{ps},Q}_{\overline{\Theta}}$ is noetherian and hence we obtain:
@@
If $\Gamma$ satisfies $\Phi_{p}$, then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}=X^{\mathrm{gen},Q}_{\overline{\Theta}}$.
@@
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}$ is represented by a finite type $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$-algebra.
@@
If $\Gamma$ satisfies $\Phi_{p}$, then $X^{\mathrm{gen},\Gamma}_{\overline{\Theta}}\sslash G^{0}\rightarrow X^{\mathrm{ps}}_{\overline{\Theta}}$ is a finite adequate homeomorphism in the sense of [1, Definition 3.3.1].
@@
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then the functor $\tilde{X}_{G}$ is representable by a quasi-Stein rigid analytic space $X_{G}$. The $\overline{L}$-points of $X_{G}$ are in canonical bijection with $G^{0}(\overline{L})$-conjugacy classes of continuous $G$-semisimple [^2] We remind the reader that a representation $\rho:\Gamma\rightarrow G(\Omega)$, where $\Omega$ is an algebraically closed field, is $G$-semisimple (or $G$-completely reducible) provided that whenever $\rho(\Gamma)$ is contained in $P(\Omega)$ for a parabolic subgroup $P$ of $G$, it is contained in $L(\Omega)$ for a Levi subgroup $L$ of $P$. We refer the reader to [2, Section 6] for more details. representations $\rho:\Gamma\rightarrow G(\overline{L})$.
@@ 1.2. What was known and tried before
Let us recall the previous related results. If $G=\operatorname{GL}_{n}$ then using the work of Emerson–Morel [10] one can work with Chenevier’s determinant laws instead of $\operatorname{GL}_{n}$-pseudocharacters. In this context 1.1 is proven by Chenevier in [6]: as part of the theory one gets a Cayley–Hamilton algebra $E$ together with a continuous representation $\rho^{\mathrm{univ}}:\Gamma\rightarrow E^{\times}$. One may then show that $\rho^{\mathrm{univ}}$ factors through $Q$ and deduce 1.1 from this, see [6, Lemma 3.8] and [14, Lemma 5.5] for details.
If $G$ is any generalised reductive group then the Cayley–Hamilton algebra is not available anymore. However, one could still try to deduce 1.2 from the $\operatorname{GL}_{n}$-case by choosing a closed immersion $\tau:G\hookrightarrow\operatorname{GL}_{n}$ and studying the map $R^{\mathrm{ps},\Gamma}_{\tau\circ\overline{\Theta}}\rightarrow R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ induced by composition with $\tau$. If the map is finite then one may deduce from the $\operatorname{GL}_{n}$-case that $R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ is noetherian. This approach was carried out by J.Q. in [16, Section 5.3] for classical groups under the assumption $p>2$ in the orthogonal cases. However, this strategy runs into difficult invariant theoretic questions in general. Since $R^{\mathrm{ps},\Gamma}_{\tau\circ\overline{\Theta}}\rightarrow R^{\mathrm{ps},\Gamma}_{\overline{\Theta}}$ is in general not surjective the authors do not see a way of deducing 1.1 from the $\operatorname{GL}_{n}$-case.
If $\Gamma$ is a finitely generated profinite group then J.Q. proves 1.2 by showing that the tangent space is finite dimensional, which is equivalent to showing that $\operatorname{Def}_{\overline{\Theta}}^{\Gamma}(k[\varepsilon])$ is finite. Let us informally explain the underlying difficulty if we drop the assumption that $\Gamma$ is finitely generated. The $G$-pseudocharacter $\Theta$ valued in the ring of dual numbers $k[\varepsilon]$ over $k$ consists of $\mathcal{O}$-algebra homomorphisms $\Theta_{n}:\mathcal{O}[G^{n}]^{G^{0}}\rightarrow\mathcal{C}(\Gamma^{n},k[\varepsilon])$ for $n\geq 1$ satisfying certain compatibility relations, where $\mathcal{C}(\Gamma^{n},k[\varepsilon])$ is the set of continuous maps. Proving finiteness assertions such as finite dimensionality of the tangent space amounts to showing that finitely many $\Theta_{n}$ determine $\Theta$ uniquely. This is feasible when the group is finitely generated, but the argument runs into trouble when this assumption is dropped. A similar problem arises in 1.6. One has to show that for an affinoid algebra $A$ a continuous $A$-valued $G$-pseudocharacter $\Theta$ takes values in some formal model of $A$. It is easy to show that each $\Theta_{n}$ takes values in $\mathcal{C}(\Gamma^{n},\mathcal{A}_{n})$ for some formal model $\mathcal{A}_{n}$ of $A$, but it is not clear that there is a formal model that works for all $\Theta_{n}$.
@@ 1.3. What we actually do
Let us explain the proof of 1.1. We first prove 1.1 for finitely generated free profinite groups $\Gamma$. We show that the map between deformation rings $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$, induced by mapping a deformation of $\overline{\rho}$ to its $G$-pseudocharacter, is injective. This key step is carried out in Section 4. Its proof uses the fact that one may describe the moduli space of $G$-pseudocharacters of a discrete finitely generated free group $F$ as a GIT-quotient of a moduli space of representations of $F$ into $G$. We show that if $\Gamma$ is the profinite completion of $F$ then the rings $R^{\mathrm{ps}}_{\overline{\Theta}}$ and $R^{\square}_{\overline{\rho}}$ are naturally isomorphic to completions of the local rings at points $\Theta|_{F}$ and $\overline{\rho}|_{F}$ in the respective moduli spaces. The assertion about injectivity boils down to commutative algebra, which is carried out in Section 3. We note that if we drop the assumption that $\Gamma$ is a free profinite group then it is not known [^3] Carl Wang-Erickson has pointed out to us that Jinyue Luo [13] has constructed an example, where the map fails to be injective. whether $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow R^{\square}_{\overline{\rho}}$ is injective, even when $G=\operatorname{GL}_{n}$, see [6, Remark 1.25].
@@
1.4. Motivation
The Hermite–Minkowski theorem implies that the Galois group $\operatorname{Gal}(F_{S}/F)$, where $F$ is a number field and $F_{S}$ is a maximal extension of $F$ unramified outside a finite set of primes $S$ of $F$, satisfies $\Phi_{p}$. It is not known if these groups are topologically finitely generated. The deformation rings of representations of $\operatorname{Gal}(F_{S}/F)$ are frequently encountered in arithmetic applications and $\Gamma=\operatorname{Gal}(F_{S}/F)$ is the main example that we have in mind in this paper.
@@ 1.5. Acknowledgements
V.P. thanks Ulrich Görtz and Marc Levine for a stimulating discussion regarding Section 3. The authors thank Gebhard Böckle, Toby Gee and Carl Wang-Erickson for their comments and the referee for their careful reading of the paper.
@@ 2. Lafforgue’s pseudocharacters
The functor $A\mapsto\mathrm{PC}^{\Gamma}_{G}(A)$ is representable by an $\mathcal{O}$-algebra [16, Theorem 3.19]. Let us write $\operatorname{Rep}^{\Gamma}_{G}(A):=\operatorname{Hom}(\Gamma,G(A))$. The functor $A\mapsto\operatorname{Rep}^{\Gamma}_{G}(A)$ is also representable by an $\mathcal{O}$-algebra. To a representation $\rho:\Gamma\to G(A)$, we associate a $G$-pseudocharacter $\Theta_{\rho}\in\mathrm{PC}^{\Gamma}_{G}(A)$ by the formula
@@ 2.1. Continuous pseudocharacters
If $\Gamma$ is a topological group, $A$ is a topological ring and, for all $n\geq 1$ the map $\Theta_{n}$ has image in the set $\mathcal{C}(\Gamma^{n},A)$ of continuous maps $\Gamma^{n}\to A$, then we say that $\Theta$ is continuous. We denote the set of continuous $G$-pseudocharacters of $\Gamma$ with values in $A$ by $\mathrm{cPC}_{G}^{\Gamma}(A)$.
@@ 2.2. Kernels
It is shown in [16, Lemma 3.11] that $\operatorname{Ker}(\Theta)$ is a normal subgroup of $\Gamma$.
@@ 2.3. Deformations of $G$-pseudocharacters
Let $\mathcal{O}$ be the ring of integers in a finite extension $L$ of $\mathbb{Q}_{p}$ with residue field $k$. Let $\Gamma$ be a profinite group and let $\overline{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(k)$. We recall a deformation problem of $\overline{\Theta}$ introduced in [16, Section 5]. Let $\mathfrak{A}_{\mathcal{O}}$ be the category of artinian local $\mathcal{O}$-algebras with residue field $k$. We define the deformation functor of $\overline{\Theta}$
@@
that sends an object $A\in\mathfrak{A}_{\mathcal{O}}$ to the set of continuous $G$-pseudocharacters $\Theta$ of $\Gamma$ over $A$ with $\Theta\otimes_{A}k=\overline{\Theta}$. It is pro-representable by a local profinite $\mathcal{O}$-algebra $R^{\mathrm{ps}}_{\overline{\Theta}}$ with residue field $k$ by [16, Theorem 5.4].
@@
Denote by $\Theta^{u}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\overline{\Theta}})$ the universal deformation of $\overline{\Theta}$. If $A$ is an $R^{\mathrm{ps}}_{\overline{\Theta}}$-algebra we will write $\Theta^{u}_{|A}$ for the specialisation of $\Theta^{u}$ along $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A$. Since $R^{\mathrm{ps}}_{\overline{\Theta}}$ is profinite and hence Hausdorff, $\operatorname{Ker}(\Theta^{u})$ is a closed normal subgroup of $\Gamma$ by 2.2.